# BiomedCoOp: Learning to Prompt for Biomedical Vision-Language Models

### v1

@@ Abstract
Recent advancements in vision-language models (VLMs), such as CLIP, have demonstrated substantial success in self-supervised representation learning for vision tasks. However, effectively adapting VLMs to downstream applications remains challenging, as their accuracy often depends on time-intensive and expertise-demanding prompt engineering, while full model fine-tuning is costly. This is particularly true for biomedical images, which, unlike natural images, typically suffer from limited annotated datasets, unintuitive image contrasts, and nuanced visual features. Recent prompt learning techniques, such as Context Optimization (CoOp) intend to tackle these issues, but still fall short in generalizability. Meanwhile, explorations in prompt learning for biomedical image analysis are still highly limited. In this work, we propose BiomedCoOp, a novel prompt learning framework that enables efficient adaptation of BiomedCLIP for accurate and highly generalizable few-shot biomedical image classification. Our approach achieves effective prompt context learning by leveraging semantic consistency with average prompt ensembles from Large Language Models (LLMs) and knowledge distillation with a statistics-based prompt selection strategy. We conducted comprehensive validation of our proposed framework on 11 medical datasets across 9 modalities and 10 organs against existing state-of-the-art methods, demonstrating significant improvements in both accuracy and generalizability. The code will be publicly available at https://github.com/HealthX-Lab/BiomedCoOp.
@@ 2 Additional Few-shot Results
Figure S1 demonstrates the performance variations of BiomedCoOp and the baseline models under different few-shot configurations ($K$=1,2,4,8, and 16). It underscores BiomedCoOp’s robustness in adapting to limited data. On the other hand, we provide the detailed few-shot evaluation results for each dataset in Table S5. Overall, BiomedCoOp is on par and regularly outperforms SOTA parameter-efficient techniques across diverse datasets.
@@ 7 Additional Hyperparameters
8 LLM Prompts Used
Central necrosis and surrounding edema in glioma tumor on MRI scan.

### v2

@@ Abstract
Recent advancements in vision-language models (VLMs), such as CLIP, have demonstrated substantial success in self-supervised representation learning for vision tasks. However, effectively adapting VLMs to downstream applications remains challenging, as their accuracy often depends on time-intensive and expertise-demanding prompt engineering, while full model fine-tuning is costly. This is particularly true for biomedical images, which, unlike natural images, typically suffer from limited annotated datasets, unintuitive image contrasts, and nuanced visual features. Recent prompt learning techniques, such as Context Optimization (CoOp) intend to tackle these issues, but still fall short in generalizability. Meanwhile, explorations in prompt learning for biomedical image analysis are still highly limited. In this work, we propose BiomedCoOp, a novel prompt learning framework that enables efficient adaptation of BiomedCLIP for accurate and highly generalizable few-shot biomedical image classification. Our approach achieves effective prompt context learning by leveraging semantic consistency with average prompt ensembles from Large Language Models (LLMs) and knowledge distillation with a statistics-based prompt selection strategy. We conducted comprehensive validation of our proposed framework on 11 medical datasets across 9 modalities and 10 organs against existing state-of-the-art methods, demonstrating significant improvements in both accuracy and generalizability. The code is publicly available at https://github.com/HealthX-Lab/BiomedCoOp.
@@ 5 Conclusion
Acknowledgment We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC).
@@ 7 Additional Few-shot Results
Figure S1 demonstrates the performance variations of BiomedCoOp and the baseline models under different few-shot configurations ($K$ = 1,2,4,8,16). It underscores BiomedCoOp’s robustness in adapting to limited data. On the other hand, we provide the detailed few-shot evaluation results for each dataset in Table S8. Overall, BiomedCoOp is on par and regularly outperforms SOTA parameter-efficient techniques across diverse datasets.
@@ 11 Selective Prompting for SCCM
12 Additional Comparisons with Other Recent Methods
We compared our method with two more recent SOTA CLIP adaptation methods (XCoOp [3] and DCPL [6]) on all datasets in Tables S2 and S3. We also compare with zero-shot methods (in blue). All methods were tuned to their optimal settings for each dataset, and 16-shot setting is used for all in Table S3. Specifically, we used an alternate version of DCPL, denoted DCPL*, which uses BiomedCLIP as the LSDM with deep multimodal prompting, whereas XCoOp directly utilizes the BiomedCLIP backbone. Our results demonstrate that our method consistently outperforms these approaches. On the other hand, average ensemble of LLM prompts improves zero-shot classification, but is suboptimal to prompt selection (Table S2).
Table S2: Avg. accuracy (%) comparison of additional SOTA methods in few-shot learning. DCPL* utilizes BiomedCLIP as the LSDM for domain knowledge while XCoOp directly utilizes the BiomedCLIP backbone. Methods in blue use zero-shot setting.
Method $K=1$ $K=2$ $K=4$ $K=8$ $K=16$
BiomedCLIP $42.05$
BiomedCLIP + Ensemble $52.27$ (+10.22(+10.22( + 10.22 from BiomedCLIP)
BiomedCLIP + Selection $53.72$ (+11.67(+11.67( + 11.67 from BiomedCLIP)
DCPL* $45.65_{8.86}$ $51.65_{8.79}$ $56.62_{7.51}$ $62.85_{8.40}$ $68.79_{4.80}$
XCoOp $52.50_{5.91}$ $55.39_{5.74}$ $60.87_{4.18}$ $66.37_{3.44}$ $71.04_{1.95}$
[…]
Table S3: Base-to-novel generalization for recent SOTA methods. DCPL* utilizes BiomedCLIP as the LSDM for domain knowledge while XCoOp directly utilizes the BiomedCLIP backbone.
Method (K=16) Base Acc. Novel Acc. HM
XCoOp 74.62% 63.19% 68.43%
DCPL* 68.70% 40.35% 50.84%
BiomedCoOp (Ours) 76.26% 73.92% 75.07%
13 Effect of LLM used
Table S4 presents an ablation study using 50 prompts from three recent LLMs across three datasets under the 4-shot evaluation. The results show that our model is robust to different text distributions, even with smaller LLMs like Gemma-2-2b, highlighting the benefit of our selection strategy in learning vigorous representations.
Table S4: Effect of different LLM choices on 4-shot accuracy (%)
Method BTMRI COVID-QU-Ex CTKidney
[…]
Gemma-2-2b 76.693.46 72.462.46 66.030.96
[…]
@@ 14 Additional Hyperparameters
15 LLM Prompts Used
Central necrosis and surrounding edema in glioma tumor on MRI scan.